INFRARED SPECTROSCOPY OF CH_{3}Cl-(ORTHO-H_{2})_{n} CLUSTERS IN SOLID PARAHYDROGEN

\underline{KAZUKI YOSHIOKA}, and DAVID T. ANDERSON, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

CH_{3}Cl doped solid parahydrogen (p-H_{2}) crystals have been studied using high-resolution FT-IR. Similar to related studies of CH_{3}F in solid p-H_{2}, multiplet structures are observed in the C-Cl stretch region, which are assigned to clustering of the CH_{3}Cl with ortho-H_{2} molecules present in the sample. In this study individual vibrational absorptions are observed for clusters in the size range of n=0-5. In contrast with the CH_{3}F studies,^{a} at low temperature (T=1.8 K) absorptions features 2.8 cm^{−1} to the blue of the main features are observed to grow in intensity. These low temperature absorption features are believed to be due to CH_{3}Cl in a new structurally distinct crystal environment that must occur by reorganization of the solid hydrogen around the CH_{3}Cl impurity at low temperature. Experimental results and analysis will be presented to support this hypothesis.

aK. Yoshioka and D. T. Anderson, J. Chem. Phys. 119, 4731 (2003).